OpenStudy (anonymous):

The equation of the straight line joining the points of intersections of x^2 + y^2 = 12 and y^2 = 4x is given by?

OpenStudy (anonymous):

1. x = -2 2. x = 2 3. y = -2 4. y = 2

OpenStudy (amistre64):

x^2 + y^2 = 12 ; replace y^2 with 4x x^2 + 4x = 12 and solve the quadratic

OpenStudy (amistre64):

most likely one solution will be extranuous, so test them out to be sure

OpenStudy (anonymous):

I got two solutions: x = -6 and x = 2. What now?

OpenStudy (amistre64):

y^2 = 4x is y^2 ever negative?

OpenStudy (amistre64):

and ... which one fits the options?